OpenStudy (anonymous):

Write the complex number in standard form. sqrt(−5)*sqrt(−10)

OpenStudy (saifoo.khan):

\[-5 \sqrt{2} \]

OpenStudy (turingtest):

\[(i\sqrt{5})(i\sqrt{10})=i^2\sqrt{50}=-5\sqrt{2}\]